Mr John bought a dozen boxes of matches. Each box contained 40 matches. One day, he used 24 match. What fraction was this of the number of matches in (a) one box (b) all twelve boxes.
Answer by Cromlix(4381)

Hi there,
Total matches in one box = 40
Mr John used 24 matches
Fraction = 24/40 cancelled down = 3/5
Total number of matches in a dozen boxes
= 12 x 40 = 480 matches
Mr John used 24 matches
Fraction = 24/480 cancelled down = 1/20.
Answers
a) 3/5
b) 1/20
Hope this helps :-)